# MERA: A Comprehensive LLM Evaluation in Russian

### v2

@@ 10 Acknowledgments
MERA is a collaborative project created in a union of industry and academia. The authors would like to express their gratitude to their partners from AI Alliance Russia, HSE University, MTS AI, and Skoltech. It is a collaboration that made an undertaking this big possible. The authors would like to extend their special thanks to Yegor Nizamov for his significant contribution in organizing the partners of the benchmark and also contractors for the creation of the website. We express our gratitude to the entire team that assist in developing the website platform and support the scoring system.
The authors are grateful to Tatiana Shavrina, who inspired this project in the first place and provided valuable ideas. The authors sincerely thank Vladislav Mikhailov and Ekaterina Taktasheva for contributing to the project and their fruitful ideas. Efforts of all the mentioned above have played a crucial role in completing this work.
@@
Appendix A Tasks Description [^16] All examples from the datasets are presented in English for transparency and are given and are for illustrative purposes only to clarify the idea of a given task. The examples are not necessarily a direct translation of specific examples from the dataset. The details about the data format and specific dataset samples are available on the project website https://mera.a-ai.ru/en/tasks.
@@ A.1.1 MathLogicQA
Math problems consist of a mathematical expression and a question about that expression. To answer the question, it is necessary to solve a linear equation or system of linear equations or perform a comparison operation. Mathematical expressions are synthetic data generated using an open-source library171717github.com/google-deepmind/mathematics_dataset using the linear_1d and linear_2d modules. The resulting generated expressions were manually rewritten by experts from mathematical language into natural Russian. Next, the experts formulated a question in natural language and the correct answer for each expression.
Experts for the correctness of the problem conditions and answers validated all examples from the dataset.
@@ A.1.2 MultiQ
MultiQ is a multi-hop QA dataset for Russian, suitable for testing general open-domain question answering, information retrieval, and reading comprehension capabilities of LLMs. The dataset is based on the dataset of the same name from the TAPE benchmark and was redesigned in the instruction format. The examples used to complement the BIG-bench were excluded from the test set.
@@ A.1.4 RCB
The Russian Commitment Bank is a corpus of naturally occurring discourse samples with a final sentence containing a clause-embedding predicate under an entailment canceling operator (question, modal, negation, antecedent of conditional). It is an instruction version of the RCB dataset from the RSG benchmark, which was additionally filtered, cleaned from the erroneous examples and augmented to ensure class balance between “entailment” and “contradiction”.
@@
ruModAr is a mathematical task from BIG-bench. The train part of the task was taken from BIG-bench repository181818github.com/google/BIG-bench/modified_arithmetic and merged into one file. The test part is new and was generated within a Python script written according to the methodology of the BIG-bench task.
@@
    102 + 435 normal-$\rightarrow$ 538
    860 + 270 normal-$\rightarrow$ 1131
    106 + 71 normal-$\rightarrow$ 178
    700 + 20 normal-$\rightarrow$ 721
    614 + 121 normal-$\rightarrow$ 736
    466 + 214 normal-$\rightarrow$
@@
ruMultiAr is a mathematical task originating from BIG-bench. The train and test parts were generated within the script from BIG-bench repository191919github.com/google/BIG-bench/multistep_arithmetic. Moreover, we added examples with division operation and, then filtered by conditions:
@@
This task tests the ability of models to solve multistep arithmetic operations (+, −, *, /). The problem is relatively simple for humans as it is solved step-by-step. Thus, the task aims to check the capability of a model to decompose complex problems into simple steps and plan actions. Moreover, sequential reasoning is one of the skills within the Fluid Intelligence ability due to the Cattell-Horn-Carroll theory of cognitive capabilities Flanagan and Dixon (2014). The purpose of ruMultiAr is to measure exactly that skill.
@@
ruOpenBookQA is a QA dataset with multiple-choice elementary-level science questions, which probe understanding of 1k+ core science facts. The original OpenBookQA Mihaylov et al. (2018) is a new kind of question-answering dataset modeled after open-book exams for assessing human understanding of a subject. It consists of 5957 multiple-choice elementary-level science questions, which probe the understanding of a small “book” of 1326 core science facts and the application of these facts to novel situations. Answering OpenBookQA questions requires additional broad common knowledge not contained in the book. The questions, by design, are answered incorrectly by both a retrieval-based algorithm and a word co-occurrence algorithm. The Russian version of the set is much smaller but covers the topics representative of the Russian language. The dataset is built with automatic translation of the original English dataset Mihaylov et al. (2018) and manual validation by a few authors; a test set was created from scratch. The set is a part of the TAPE benchmark that was redesigned to an instruction format and filtered. The samples that are part of the BIG-bench set were excluded.
@@
Turing-test Interview Emulation (ruTiE) is a simulation of the Turing test [^20] https://plato.stanford.edu/entries/turing-test/ in Russian. The dataset was collected manually by markers and then validated. The first version of the dataset consists of only one long dialogue of length 430 for the training public set and one dialogue of length 430 for the test set. The dataset imitates a coherent dialogue with the subject, where the subject is asked questions on various topics, covering multiple categories (sentiment, intent, style, humor, irony, facts, profanity, text metrics, language structure, topic modeling, multilanguage, algorithmic transformation) of different aspects of human cognition. The subject needs to choose which of the two answer options is correct. ruTiE questions imply that the subject (model) fully remembers the context of the dialogue [^21] The dialogue context is composed of the previous questions and the answer options chosen by the subject in prior steps. For context questions, there is no information about all possible answer options. and may have a reference to the previous parts. Another peculiarity of the dataset is that the answers are not binary (correct vs. not correct). One should process both answers to give the correct response.
@@
ruWorldTree is a QA dataset with multiple-choice elementary-level science questions that evaluate the understanding of core science facts. The set is created based on the original English WorldTree dataset Jansen et al. (2018) that provides a corpus of explanation graphs for elementary science questions. The data includes the corpus of factoid utterances of various kinds, complex factoid questions, and a corresponding causal chain of facts from the corpus, resulting in a correct answer. The set is a part of the TAPE benchmark that was redesigned to an instruction format, verified, and cleaned from the erroneous samples and the BIG-bench samples.
@@ A.2.1 BPS
The Balanced Parentheses Sequence is an algorithmic task originating from BIG-bench. The primary purpose of this task is to measure language models’ ability to learn CS algorithmic concepts like stacks, recursion, or dynamic programming. Each subtask contains a parentheses sequence. The model’s goal is to predict whether the sequence is balanced or not correctly. The parentheses sequences of the length 2, 4, 8, 12, and 20 were generated for the train and test sets within a Python script.
@@ A.2.2 CheGeKa
CheGeKa is a Jeopardy!-like242424www.jeopardy.com Russian QA dataset collected from the official Russian quiz database ChGK Mikhalkova and Khlyupin (2022) and belongs to the open-domain question-answering group of tasks. The dataset is based on the corresponding dataset from the TAPE benchmark Taktasheva et al. (2022). The examples used to complement the BIG-bench Srivastava et al. (2023) were excluded from the test set.
@@ A.2.3 LCS
The Longest Common Subsequence (LCS) is an algorithmic task originating from BIG-bench. This problem consists of pairs of strings as an input, and language models are expected to correctly predict the length of the longest common subsequence between the strings. Sequences of different lengths were generated with a Python script for train and test sets.
@@
ruHumanEval is the Russian counterpart of the HumanEval dataset Chen et al. (2021), assessing models’ abilities to generate solutions for straightforward programming problems on Python. The training part of the dataset contains the translated into Russian and manually verified tasks of the original dataset [^25] https://huggingface.co/datasets/openai_humaneval including the test cases, which was taken from Liu et al. (2023a) (10 test cases per task). The test part is created from scratch by assembling various programming tasks of the same difficulty level as the training part and manually writing the test cases and documentation strings. All tasks were verified to ensure no repetitions of the training samples. This task evaluates the functional correctness of code generation by providing input information, including a textual function description (docstring) and examples of expected results for different test cases.
@@
ruMMLU is created based on the original MMLU dataset Hendrycks et al. (2021b) and follows its methodology. The dataset is designed to evaluate elementary knowledge and expertise in various domains acquired by a model during pre-training.
The training part of the dataset is created from the translated into Russian and additionally verified and filtered (via TagMe [^26] https://tagme.sberdevices.ru platform) tasks of the original dataset [^27] https://huggingface.co/datasets/cais/mmlu. The test part is new and was created as a part of the MERA project. This part contains tasks that cover the exact domains and subdomains as the train one while keeping them all balanced and including more Russian historical and cultural facts.
@@ A.2.6 USE
The dataset comprises tasks from the Unified State Exam [^28] https://fipi.ru/ege (USE) for graduates of Russian schools. The exam consists of 27 question types: 26 test-type tasks and writing an essay based on a fiction text. Each task is designed to measure proficiency in specific domains of the Russian language, such as spelling, orthoepy, grammar, punctuation, stylistics, semantics, and text interpretation. The content of the exam may vary depending on the year. The benchmark included tasks and assessment criteria for the USE 2019.
@@
  • text: open-question tasks (tasks 2, 4-7, 13, 14, 24);
  • multiple_choice: tasks that require to choose one or more correct answers from the given answer options (tasks 1, 3, 8-12, 15-23, 25) and are divided into three subtypes: based_on_text consist of text, text-based question and answer options, options_within_text — text and answer options in the text, independent_options — question and answer options;
@@ A.3 Diagnostic Datasets
ruDetox diagnostic is a part of ruDetox dataset Dementieva et al. (2022), a parallel corpus for text detoxification. For this task we took the publicly available dev split of the dataset292929github.com/s-nlp/russe_detox_2022/dev.tsv. The task is to rewrite the original toxic comment in a non-toxic way. Thus, it can be viewed as a Textual Style Transfer problem Dementieva et al. (2021); Dale et al. (2021); Logacheva et al. (2022), where the goal is to reformulate the sentence in a non-toxic style, preserving original meaning and fluency.
@@
ruEthics is an ethical diagnostic dataset aimed at assessing how LLMs perceive the fundamental concepts of ethics and how these concepts relate to the five fundamental ethical norms from Hendrycks et al. (2021a): virtue, law, morality, justice, and utilitarianism. The dataset is based on data from ethical datasets (Ethics1) and Ethics2) from the TAPE benchmark, which were revised and relabelled for the current benchmark.
@@
The evaluation process allows us to obtain the “overall ethical portrait of the model”, i.e., how the most general concepts related to ethics are decomposed for the model according to these 5 criteria.
As for the dataset creation, the texts from the original Ethics datasets (only the train part) were filtered based on the inconsistency of annotators’ responses, the length of the texts, and their style and augmented with 12 poetic texts. Then, the actants of each text were annotated via the ABC platform with 3 votes per each text, and for each text, five random actants were selected and organized in the ordered pairs (no more than 20 pairs per text). Finally, for the human benchmark, the resulting samples were sent for annotation via the ABC platform with the questions for the language models.
@@
ruHateSpeech dataset is inspired by the English HateSpeech dataset de Gibert et al. (2018) and designed to analyze a model’s proficiency in identifying negative statements addressed to specific social groups (men, women, migrants, nationality, LGBT, other) through binary-choice questions. Following the methodology and main traits of the original dataset, the replicas and replies were collected and verified through a multi-stage annotation on Toloka to ensure that the replies are fluent, coherent, related to the initial replica, and at least one reply to the same replica is toxic towards a particular social group. After annotation, only those quadruples (one replica, two replies, and one target group) that were chosen by 3 annotators were manually verified to prevent possible leakages and then formed the ruHateSpeech dataset.
Each example consists of one replica and two possible replies to it. Replies are constructed in such a way that one reply is always toxic towards the given target group while the other one is usually also toxic but towards the other social group. The replies are constructed so that the presence of obscene, swearing, or abusive vocabulary does not indicate the right option, as the other option also contains such constructions. The diagnostic dataset facilitates an initial assessment, establishing ethical guidelines for evaluating the model’s proficiency in recognizing various biases and stereotypes.
@@
ruHHH (Helpful, Honest & Harmless Alignment) is a robust LLM evaluation tool employing binary-choice tasks that were created through manual translation of the original BIG-Bench dataset tasks [^30] https://github.com/google/BIG-bench/tree/main/bigbench/benchmark_tasks/hhh_alignment Askell et al. (2021) and additional verification to ensure the texts reflect the Russian cultural traits. Despite the mentioned changes ruHHH is comparable with the original English dataset.
@@
Appendix B Motivation for the Selection of the Number of Few-shot Examples

### v3

@@ 10 Acknowledgments
MERA is a collaborative project created in a union of industry and academia. The authors would like to express their gratitude to their partners from AI Alliance Russia, HSE University, MTS AI, and the Center for Artificial Intelligence Technology. It is a collaboration that made an undertaking this big possible. The authors would like to extend their special thanks to Yegor Nizamov for his significant contribution in organizing the benchmark partners and contractors for creating the website. We express our gratitude to the entire team that assists in developing the website platform and supports the scoring system.
The authors are grateful to Tatiana Shavrina, who inspired this project in the first place and provided valuable ideas. The authors sincerely thank Vladislav Mikhailov and Ekaterina Taktasheva for contributing to the project and their fruitful ideas. We express our gratitude to Daria Latortseva, who contributed a lot to the ruMMLU dataset creation. Efforts of all the abovementioned have played a crucial role in completing this work.
@@
Appendix A Tasks Description [^17] All examples from the datasets are provided in English for illustrative purposes to clarify the concept of a given task. The examples are not necessarily a direct translation of specific examples from the dataset. The details about the data format and specific dataset samples are available on the project website https://mera.a-ai.ru/en/tasks.
@@ A.1.1 MathLogicQA
Math problems consist of a mathematical expression and a question about that expression. To answer the question, it is necessary to solve a linear equation or system of linear equations or perform a comparison operation. Mathematical expressions are synthetic data generated using an open-source library [^18] https://github.com/google-deepmind/mathematics_dataset using the linear_1d and linear_2d modules. The resulting generated expressions were manually rewritten by experts from mathematical language into natural Russian. Next, the experts formulated a question in natural language and the correct answer for each expression.
All examples were validated via the Toloka annotation platform. As a result of validation, the final test sample included examples with the entire expert agreement. The training set included the remaining examples with agreement above 60%. See Tab. 7 for more details.
@@ A.1.2 MultiQ
MultiQ is a multi-hop QA dataset for Russian, suitable for testing general open-domain question answering, information retrieval, and reading comprehension capabilities of LLMs. The dataset is based on the dataset of the same name from the TAPE benchmark Taktasheva et al. (2022) and was redesigned in the instruction format. The examples used to complement the BIG-bench were excluded from the test set.
@@ A.1.4 RCB
The Russian Commitment Bank is a corpus of naturally occurring discourse samples with a final sentence containing a clause-embedding predicate under an entailment canceling operator (question, modal, negation, antecedent of conditional). It is an instruction version of the RCB dataset from the RSG benchmark, which was additionally filtered, cleaned from the erroneous examples, and augmented to ensure a class balance between “entailment” and “contradiction”.
@@
ruModAr is a mathematical task from BIG-bench. The public test part of the task was taken from BIG-bench repository [^19] https://github.com/google/BIG-bench/modified_arithmetic and merged into one file. The test part is new and was generated within a Python script written according to the methodology of the BIG-bench task.
@@
    102 + 435 $\rightarrow$ 538
    860 + 270 $\rightarrow$ 1131
    106 + 71 $\rightarrow$ 178
    700 + 20 $\rightarrow$ 721
    614 + 121 $\rightarrow$ 736
    466 + 214 $\rightarrow$
@@
ruMultiAr is a mathematical task originating from BIG-bench. The train and test parts were generated within the script from BIG-bench repository [^20] https://github.com/google/BIG-bench/multistep_arithmetic. Moreover, we added examples with division operation and, then filtered by conditions:
@@
This task tests the ability of models to solve multistep arithmetic operations (+, −, *, /). The problem is relatively simple for humans as it is solved step-by-step. Thus, the task aims to check a model’s capability to decompose complex problems into simple steps and plan actions. Moreover, sequential reasoning is one of the skills within the Fluid Intelligence ability due to the Cattell-Horn-Carroll theory of cognitive capabilities Flanagan and Dixon (2014). The purpose of ruMultiAr is to measure exactly that skill.
@@
ruOpenBookQA is a QA dataset with multiple-choice elementary-level science questions, which probe understanding of 1k+ core science facts. The original OpenBookQA Mihaylov et al. (2018) is a new kind of question-answering dataset modeled after open-book exams for assessing human understanding of a subject. It consists of 5957 multiple-choice elementary-level science questions, which probe the understanding of a small “book” of 1326 core science facts and the application of these facts to novel situations. Answering OpenBookQA questions requires additional broad common knowledge not contained in the book. The questions, by design, are answered incorrectly by both a retrieval-based algorithm and a word co-occurrence algorithm. The Russian version of the set is much smaller but covers the topics representative of the Russian language. The dataset is built with automatic translation of the original English dataset Mihaylov et al. (2018) and manual validation by the authors; a test set was created from scratch. The set is a part of the TAPE benchmark that was redesigned to an instruction format and filtered. The samples that are part of the BIG-bench set were excluded.
@@
Turing-test Interview Emulation (ruTiE) is a simulation of the Turing test [^21] https://plato.stanford.edu/entries/turing-test in Russian. The dataset was collected manually and then validated by annotators. The first version of the dataset consists of only one long dialogue of length 430 for the training set and one dialogue of length 430 for the test set. The dataset imitates a coherent dialogue with the subject, where the subject is asked questions on various topics, covering multiple categories (sentiment, intent, style, humor, irony, facts, profanity, text metrics, language structure, topic modeling, multilanguage, algorithmic transformation) of different aspects of human cognition. The subject needs to choose which of the two answer options is correct. ruTiE questions imply that the subject (model) fully remembers the context of the dialogue [^22] The dialogue context is composed of the previous questions and the answer options chosen by the subject in prior steps. There is no information about all possible answer options for context questions. and may have a reference to the previous parts. Another peculiarity of the dataset is that the answers are not binary (correct vs. incorrect). One should process both answers to give the correct response.
@@
ruWorldTree is a QA dataset with multiple-choice elementary-level science questions that evaluate the understanding of core science facts. The set is created based on the original English WorldTree dataset Jansen et al. (2018) that provides a corpus of explanation graphs for elementary science questions. The data includes the corpus of factoid utterances of various kinds, complex factoid questions, and a corresponding causal chain of facts from the corpus, resulting in a correct answer. The set is part of the TAPE benchmark redesigned to an instruction format, verified, and cleaned from the erroneous and BIG-bench samples.
@@ A.2.1 BPS
The Balanced Parentheses Sequence is an algorithmic task originating from BIG-bench. This task’s primary purpose is to measure language models’ ability to learn CS algorithmic concepts like stacks, recursion, or dynamic programming. Each subtask contains a parentheses sequence. The model’s goal is to predict whether the sequence is balanced or not correctly. For the train and test sets, parentheses sequences of lengths 2, 4, 8, 12, and 20 were generated using a Python script.
@@ A.2.2 CheGeKa
CheGeKa is a Jeopardy!-like [^25] https://www.jeopardy.com Russian QA dataset collected from the official Russian quiz database ChGK Mikhalkova and Khlyupin (2022) and belongs to the open-domain question-answering group of tasks. The dataset is based on the corresponding dataset from the TAPE benchmark Taktasheva et al. (2022). The examples used to complement the BIG-bench Srivastava et al. (2023) were excluded from the test set.
@@ A.2.3 LCS
The Longest Common Subsequence (LCS) is an algorithmic task originating from BIG-bench. This problem consists of pairs of strings as an input, and language models are expected to correctly predict the length of the longest common subsequence between the strings. The latter varies from 0 to 9. Thus, the task can be regarded as a ten-class classification problem.
The public test part of the task was taken from BIG-bench repository [^26] https://github.com/google/BIG-bench/tree/main/bigbench/benchmark_tasks/cs_algorithms/lcs.
For the test set sequences of different lengths were generated using a Python script.
@@
ruHumanEval is the Russian counterpart of the HumanEval dataset Chen et al. (2021), assessing models’ abilities to generate solutions for straightforward programming problems on Python. The public test of the dataset contains the translated into Russian and manually verified tasks of the original dataset [^27] https://huggingface.co/datasets/openai_humaneval including the test cases, which was taken from Liu et al. (2023a) (10 test cases per task). The test part is created from scratch by assembling various programming tasks of the same difficulty level as the public test part and manually writing the test cases and documentation strings. All tasks were verified to ensure no repetitions of the public test samples. This task evaluates the functional correctness of code generation by providing input information, including a textual function description (docstring) and examples of expected results for different test cases.
@@
ruMMLU is created based on the original MMLU dataset Hendrycks et al. (2021b) and follows its methodology. The dataset is designed to evaluate expertise in various domains acquired by a model during pre-training.
The public test part of the dataset was created from the translated into Russian and additionally filtered (via the TagMe platform) tasks of the original dataset [^28] https://huggingface.co/datasets/cais/mmlu. During filtration on a platform, about 220 unique annotators labeled the text translations and checked the translation’s correctness, with an overlap equal to 5. The aggregation strategy of labeling was handled with the GLAD algorithm Whitehill et al. (2009) with the threshold equal to 0 to maximize the number of labels agreed between 5 answers from the annotators. After that, approximately 5,000 tasks, filtered out as poorly translated according to the annotators, were correctly handwritten by experts.
The closed test part was collected manually by experts as a part of the MERA project following MMLU methodology. This part contains tasks that cover the exact domains and subdomains as the public test one while keeping them all balanced and including more Russian historical and cultural facts.
@@ A.2.6 USE
The dataset comprises tasks from the Unified State Exam [^29] https://fipi.ru/ege (USE) for graduates of Russian schools. The exam consists of 27 questions: 26 test-type tasks and writing an essay based on a fiction text. Each task is designed to measure proficiency in specific domains of the Russian language, such as spelling, orthoepy, grammar, punctuation, stylistics, semantics, and text interpretation. The content of the exam may vary depending on the year. The benchmark included tasks and assessment criteria for the USE 2019.
@@
  • text: open-question tasks (tasks 2, 4–7, 13, 14, 24);
  • multiple_choice: tasks that require to choose one or more correct answers from the given answer options (tasks 1, 3, 8–12, 15–23, 25) and are divided into three subtypes: based_on_text consist of text, text-based question and answer options, options_within_text — text and answer options in the text, independent_options — question and answer options;
@@
All tasks are rated in complete concordance with the official USE assessment guide. The grading system is as follows:
  • For correct completion of tasks 1–15 and 17–25, the examinee receives 1 point. For an incorrect answer or lack of an answer, the examinee receives 0 points.
  • For completing task 16, the examinee receives from 0 to 2 points. The examinee receives 2 points if all numbers are correct. One point is given if one of the numbers in the answer is incorrect or one of the numbers in the answer is missing. In all other cases, 0 points are given.
  • For completing task 26, the examinee receives from 0 to 4 points. The examinee receives 4 points if all numbers are correct. For each correctly indicated number, the examinee receives 1 point.
The final metric is the Grade norm score, the average normalized primary score across all versions. The primary score is the sum of points for all exam tasks.
For the text and multiple_choice tasks from the test sample, for which the answer is a string containing several words or a string containing a sequence of numbers, all possible combinations of these words and numbers are used when calculating metrics. Only one answer combination is presented for these tasks from the train and dev sets.
@@ A.3 Diagnostic Datasets
ruDetox diagnostic is a part of ruDetox dataset Dementieva et al. (2022), a parallel corpus for text detoxification. For this task we took the publicly available dev split of the dataset [^30] https://github.com/s-nlp/russe_detox_2022/dev.tsv. The task is to rewrite the original toxic comment in a non-toxic way. Thus, it can be viewed as a Textual Style Transfer problem Dementieva et al. (2021); Dale et al. (2021); Logacheva et al. (2022), where the goal is to reformulate the sentence in a non-toxic style, preserving original meaning and fluency.
@@
ruEthics is an ethical diagnostic dataset aimed at assessing how LLMs perceive the fundamental concepts of ethics and how these concepts relate to the five fundamental ethical norms from Hendrycks et al. (2021a): virtue, law, morality, justice, and utilitarianism. The dataset is based on data from ethical datasets (Ethics1) and Ethics2) from the TAPE benchmark, which was revised and relabelled for the current benchmark.
@@
The evaluation process allows us to obtain the “overall ethical portrait of the model”, i.e., how the model’s most general concepts related to ethics are decomposed according to these 5 criteria.
As for the dataset creation, the texts from the original Ethics datasets (only the train part) were filtered based on the inconsistency of annotators’ responses, the length of the texts, and their style and augmented with 12 poetic texts. The data annotation process for the dataset included parsing the actants from preprocessed texts via the ABC platform. For each text, 3 different annotators proposed their variants of the actants involved in actions from the text. The results have been aggregated to form the proposed list of actants for each text. The expert verified and filtered these lists to form the resulting arrays of actants for each dataset text. Finally, for the human benchmark, the resulting samples were sent for annotation via the ABC platform with the questions for the language models. See Tab. 7 for more annotation details.
@@
ruHateSpeech dataset is inspired by the English HateSpeech dataset de Gibert et al. (2018) and designed to analyze a model’s proficiency in identifying negative statements addressed to specific social groups (men, women, migrants, nationality, LGBT, other) through binary-choice questions.
Each example consists of one replica and two possible replies to it. Replies are constructed in such a way that one reply is always toxic towards the given target group while the other one is usually also toxic but towards a different social group. The replies are constructed so that the presence of obscene, swearing, or abusive vocabulary does not indicate the right option, as the other option also contains such constructions. The diagnostic dataset facilitates an initial assessment, establishing ethical guidelines for evaluating the model’s proficiency in recognizing various biases and stereotypes.
Following the methodology and main traits of the original dataset, the replicas and replies were collected and verified through a multi-stage annotation on Toloka to ensure that the replies are fluent, coherent, related to the initial replica, and at least one reply to the same replica is toxic towards a particular social group. See Tab. 7 for data annotation details.
@@
ruHHH (Helpful, Honest & Harmless Alignment) is a robust LLM evaluation tool employing binary-choice tasks that were created through manual translation of the original BIG-Bench dataset tasks [^31] https://github.com/google/BIG-bench/tree/main/bigbench/benchmark_tasks/hhh_alignment Askell et al. (2021) and additional verification to ensure the texts reflect the Russian cultural traits. Despite the mentioned changes, ruHHH is comparable with the original English dataset.
@@
Appendix B Incorporating Culture-specific Context in MERA Benchmark
A comprehensive benchmark should evaluate culture-specific and local-specific aspects of the culture and the language it is aimed at. Understanding this, we ensured the presence of Russian-specific aspects in MERA:
  • Several tasks in MERA are based on problems of Russian origin or were created specifically for the MERA benchmark. The USE dataset contains questions from the Russian Language United States Exam and naturally reflects the Russian-language specifics. CheGeKa contains questions from the Russian game “What? Where? When?” (or “Chto? Gde? Kogda?” in transliteration), which is a Russian analog of the Jeopardy! Game. This game resembles the Jeopardy! Show only in its idea and format, while its questions usually include cultural and historical specifics of Russia (e.g., questions about famous Russian artists, poets, or historical figures). ruTiE is a new dataset developed and created specifically for the benchmark. During its creation, we ensured that it included questions requiring knowledge of Russian cultural and historical specifics.
  • Other tasks (e.g., RCB and PARus) were created using originally Russian textual data such as news, articles, and book corpora and, thus, contain texts mentioning cultural, social, and political aspects.
  • For some datasets, we translated only the public test set while creating the entirely new closed test part (the only part used for the evaluation) from scratch (e.g. ruMMLU) specifically for the MERA benchmark. In such cases, we included more questions about Russian historical and cultural facts while preserving the original domain and subdomain distribution.
  • There are datasets (e.g., ruHHH, ruWorldTree, RWSD) where we conducted a cultural adaptation, which included replacing cultural and historical concepts with Russian-related ones. For ruHHH, an example of such adaptation is presented in the dataset description in Sec. A.3.4.
It should also be noted that MERA includes tasks that evaluate math and computer code skills. These tasks are language-agnostic and, thus, do not require any language adaptation.
Appendix C Motivation for Metric Selection
We use a set of metrics for the evaluation at the benchmark tasks. The description of the metrics can be found in Sec. 4.2, and the metric for each task is specified in Tab. 2. For the datasets that were adapted, translated, or based on some other dataset, we mostly used metrics for scoring the original task. Namely:
  • for LCS, BPS, and ruHHH we used metrics from the corresponding BIG-bench tasks Srivastava et al. (2023);
  • for ruModAr, ruMultiAr, and SimpleAr inline with the BIG-bench approach, we measure the percentage of the correct answers and compare model predictions with golden answers using EM;
  • for PARus, RCB, and RWSD we followed RSG methodology Shavrina et al. (2020b);
  • for MultiQ, ruOpenBookQA, ruWorldTree, CheGeKa we used the same metrics as in TAPE Taktasheva et al. (2022);
  • for ruMMLU we adopted the original MMLU Hendrycks et al. (2021b) approach scoring it with Accuracy;
  • for ruHateSpeech we adapted the methodology of the English HateSpeech dataset de Gibert et al. (2018);
  • for ruHumanEval we repeated the scoring procedure for the original HumanEval dataset Chen et al. (2021);
  • for ruDetox we used the Joint score employed for the original task Logacheva et al. (2022).
As for the other tasks, we selected the metric based on the task formulation, task answer type, and the task-specific details:
  • we scored ruTiE and MathLogicQA using accuracy as the answers in the datasets are balanced, and it is a standard benchmark metric for binary classification tasks;
  • for ruEthics we adopted the methodology of the GLUE Wang et al. (2018) diagnostic dataset extending it to the 5 ethical criteria. The motivation for this was the class imbalance and the absence of the actual golden answer (see Sec. A.3.2 for the task details);
  • for USE, we use Grade norm score, the average normalized primary score across all versions. The primary score is calculated according to the official USE assessment guide [^32] https://fipi.ru/ege (see Sec. A.2.6 for details).
Appendix D Motivation for the Selection of the Number of Few-shot Examples